OpenStudy (anonymous):

if i have a symmetrical set up of charges within the center of a cube, how do i calculate the E flux through one side of the cube?

OpenStudy (ghazi):

electric flux through one side of the cube will be ...as we know by gauss law flux= \[\frac{ q }{ \epsilon }\] therefore for one side it will be =\[\frac{ q }{ 6 \epsilon }\]

OpenStudy (ghazi):

since total flux= q/ epsilon therefore through one side it'll be = q/6 epsilon...cube has got 6 faces

OpenStudy (anonymous):

Thanks man really appreciate it!

OpenStudy (ghazi):

:)